OpenStudy (anonymous):

Simplify each product 21/81 x 63/105

OpenStudy (anonymous):

21/81 X 63/105= 1/9 x 7/5 = 7/45 need to find lowest product then multiply

jhonyy9 (jhonyy9):

21 63 7*3 7*9 1 7 7 --- * ---- = ---- * ------ = --- * --- = --- 81 105 9*9 7*3*5 9 5 45

OpenStudy (anonymous):

y=x-v/b